Down Under Boomerang, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2.52 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,020,000 in annual sales, with costs of $715,000. The tax rate is 30 percent and the required return is 16 percent. What is the project’s NPV? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) What is NPV?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,520,000/3)=$840,000/year

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(2,020,000-715000)(1-0.3)+(0.3*840,000)

=$1165500

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=$1165500[1-(1.16)^-3]/0.16

=$1165500*2.24588954

=$2,617,584.26

NPV=Present value of inflows-Present value of outflows

=$2,617,584.26-$2,520,000

=$97,584.26(Approx).